OpenStudy (anonymous):

How do i simplify the square root of 45?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\sqrt{45}\\ =\sqrt{9\times 5}\\ =\sqrt{9}\times\sqrt{ 5}\\ =3\times\sqrt{ 5}\\ =3\sqrt{5}}\end{align}\)

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

pull out the perfect squares in the square root of 45 so for example you can do the square root of 9 and square root of 5 your final answer would be 3 square root 5